OpenStudy (anonymous):

Express 1−sqrt*3i in polar form R(cosθ + i sinθ) with −π < θ ≤ π.

ganeshie8 (ganeshie8):

R = magnitude = \(\sqrt{1^2+(-\sqrt{3})^2} = \sqrt{1+3} = 2\)

ganeshie8 (ganeshie8):

\(\theta = \arctan (\dfrac{-\sqrt{3}}{1}) = \arctan(-\sqrt{3}) = ?\)

ganeshie8 (ganeshie8):

plugin \(R, \theta\) values in the given polar form ^

OpenStudy (anonymous):

i havent done polar coordinates yey, so i dont get the question